A salesperson has a 20 percent chance of making a sale to any customer who is called. If you make 6 calls, what is the probability that

A) Three calls are made

B) Probability of at least 1 sale is made.

C) Find the mean and standard deviation.
Answer #1

A) Three calls are made

P(x = 3) = 6C3*0.2^3*0.8^3

P(x=3) = 20*0.008*0.512 = 0.0819

B) Probability of at least 1 sale is made.

P(1<= x <=6) = 1 - P(x = 0) = 1 - 0.8^6 =1-0.2621 = 0.7379

C) Find the mean and standard deviation.

mean = np = 6*0.2 = 1.2

stdev = sqrt(npq) = sqrt(1.2*0.8) = 0.979